OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (anonymous):

Point-slope form: y - y1 = m(x - x1) y -(-2) = 2(x - 3) y + 2 = 2x - 6 -2 y = 2x - 8 Hope this helps!

OpenStudy (anonymous):

Thank you very much, Jazy.